OpenStudy (anonymous):

write an equation for the line that is parallel to the given line and passes through the given point Y=3x+7; (2, 10) A. Y=3x+11 B. y= 3x+4 C. Y=-3x+4 D. Y=(1/3)x+4

OpenStudy (anonymous):

The slope of the parallel line will have the same slope as the given line.

OpenStudy (anonymous):

Use the formula y = mx+b....and the point (2,10) to help finish the equation

OpenStudy (anonymous):

hint they have the same slope so plug them in to \[y-y_1 = m (x-x_1)\] then put it into \[y = mx+b \] form. the M is the slope, the b is the y -intercept

OpenStudy (anonymous):

Paralell means the slope are the same (meaning that the slope = 3), so choice D is eliminated. A. Y = 3x + 11 plug in (2, 10) 10 = 3(2) + 11 10 = 17 false. B. Y = 3x + 4 plug in (2, 10) 10 = 3(2) + 4 10 = 10 So the answer is B.